# Surfel-based Gaussian Inverse Rendering for Fast and Relightable Dynamic Human Reconstruction from Monocular Videos

### v2

@@ 3.4 Reconstruct animatable PBR-aware 2DGS from monocular videos
Our training process consists of two stages. In the first stage, we apply an Image Reconstruction loss to capture the rough shape of the dynamic humans. In the second stage, we use Physically-Based Rendering (PBR) to estimate PBR materials and achieve accurate surface reconstruction. We add additional regularization terms to obtain a suitable solution for such a highly ambiguous problem. Specifically, given $M$ frames from a monocular video and the estimated poses of the human $\{(\mathbf{I}_{1},\boldsymbol{\theta}_{1}),\dots,(\mathbf{I}_{M},\boldsymbol{\theta}_{M})\}$ we optimize $\mathbf{G},\widehat{\mathbf{W}},\widetilde{\mathbf{B}},\widetilde{\mathbf{W}}$ with the ground-truth images.
Training Stage I – Image Reconstruction. In the first stage, we aim to reconstruct the rough shape for clothed avatars with view-dependent color represented as spherical harmonics. We apply the image reconstruction loss, regularization term on these Gaussian parameters, and the normal-depth consistency loss proposed by [33]. Denote the rendered image as $\hat{\mathbf{I}}$, the training loss at this stage is:
$L_{\text{RF}}=\ell_{1}(\hat{\mathbf{I}},\mathbf{I})+\lambda_{\text{SSIM}}L_{\text{SSIM}}(\hat{\mathbf{I}},\mathbf{I})+L_{\text{reg}}+\lambda_{\text{NC}}L_{\text{NC}}$ (15)
we only render the foreground avatar and apply the $L_{\text{RF}}$ loss on the foreground avatar and do not model the background. Due to Gaussians’ discrete representation nature, without any regularization term, the reconstructed cloth avatars may be overfitting to the sparse 2D observation. We therefore follow previous works [40] and regularize the standard variation $\sigma_{k}$ of Gaussian attributes in the KNN neighborhood of each Gaussian $\mathbf{G}_{k}$, which leads to:
$L^{\sigma}_{k}=\sum_{\text{attr}_{k}\in\{\mathbf{R}_{k},\mathbf{S}_{k},o_{k},\mathbf{c}_{k},\Delta\mathbf{W}_{k}\}}\lambda_{\text{attr}_{k}}\sigma(\text{attr}_{k})$ (16)
where $\sigma_{\text{attr}_{k}}$ denotes the standrd variation of the Gaussian attributes in the KNN neighborhood. Furthermore, we encourage learnable blending weight, and the latent bones for cloth animation for each Gaussian should also be small, thus alleviating the overfitting for the cloth deformation. We follow the previous work [35] to store the learnable blending weight in a 3D voxel and apply a tri-linear interpolation to query the blending weight $\Delta W_{k}$ for each Gaussian. We therefore regularize the learnable blending weight to be small for both the SMPL bones and latent bones with:
$L_{k}^{\text{norm}}=\lambda_{\widehat{\mathbf{W}}}\|\Delta\mathbf{W}_{k}\|_{2}+\lambda_{\widetilde{\mathbf{W}}}\|\widetilde{\mathbf{W}}_{k}\|_{2}$ (17)
@@
$L_{\text{reg}}=\frac{1}{N_{\text{gs}}}\sum_{k=1}^{N_{\text{gs}}}L_{k}^{\sigma}+L_{k}^{\text{norm}}$ (18)
To ensure the orientation of each 2D Gaussians locally aligned with the actual surface and suitable for physically based rendering. We follow 2DGS [33] and align the splat’s normal with the gradients of the depth maps as follows:
$L_{\text{NC}}=\sum_{i}w_{i}(1-\mathbf{n}_{i}^{T}\mathbf{N})$ (19)
where $w_{i}=o_{i}\mathcal{G}_{i}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))$ denotes the blending weight of the $i$-th intersection, $\mathbf{n}_{i}$ is the normal of the splat facing the camera, $\mathbf{N}$ is normal derived from depth map, In contrast to the $\ell_{1}$ loss proposed by previous works [47, 49], we find this normal-consistency loss can achieve more smooth surface and consistent orientation for each gaussian along the ray, as validated in our ablation study.
Training Stage II – PBR optimization. At the PBR stage, we first apply the image reconstruction loss on the PBR image. Denote the rendered image of PBR as $\mathbf{\hat{I}}_{\text{pbr}}(\mathbf{a},m,r,\mathbf{n})$, which is computed from albedo, materials, and splat normal. We apply:
@@
$\lambda_{\text{NC}}L_{\text{NC}^{*}}+\lambda_{\text{white}}L_{\text{white}}$
Here the $L_{\text{NC}^{*}}$ is slightly different from the $L_{\text{NC}}$ used in the previous stage. With original normal-consistency loss from the previous stage, we can not reconstruct totally accurate geometry, especially under the extreme high light conditions as shown in Fig 3 and original 2DGS [33] paper. At the PBR stage, the appearance of Gaussians are correlated with the splat normal. Thus we can further improve the geometry during the optimization. Therefore, we can not freeze the Gaussian shape during the PBR optimization stage as proposed by GS-IR [47]. However, if we directly optimize with Equation 19, the splat normal still needs to be consistent with the inaccurate geometry and can not improve the geometry of clothed avatars. If we discard this normal consistency constraint, the splat normal will be totally messy during the optimization. Therefore, we propose a progressive optimization strategy to achieve both accurate geometry and consistent splat normal with following:
@@
$+\lambda_{\text{NC}_{2}}\sum_{i}w_{i}(1-\mathbf{n}_{i}^{T}\mathbf{N})$
where $\text{sg}(\cdot)$ denotes the stop gradient. At the initial stage, we set $\lambda_{\text{NC}_{2}}=0,\lambda_{\text{NC}_{1}}=1$, which means we utilize the splat normal to optimize the geometry $\mathbf{N}$. Then we set $\lambda_{\text{NC}_{1}}=0,\lambda_{\text{NC}_{2}}=1$ to force the splat normal to be aligned with the fixed geometry. We visualize the splat normal and depth derived normal during the optimization stage at Fig 3. To overcome the ambiguity between the lighting and PBR materials and obtain a plausible solution for PBR materials, we follow  [48] to utilize a smoothness loss which forces the materials and albedo to be smooth at the planar surfaces where the image gradients are absent. This can be written as:
@@
$+\frac{1}{N_{\text{pixel}}}\sum_{i,j}|\partial_{\mathbf{y}}\mathbf{a}_{ij}|e^{-\|\partial_{\mathbf{y}}\mathbf{I}_{ij}\|}$
where $N_{\text{pixel}}$ denotes the numbe of pixels for the image. I represents the ground-truth images, $\partial_{\mathbf{x}},\partial_{\mathbf{y}}$ represents the gradients of an image or material map in $\mathbf{x}$ and $\mathbf{y}$ directions. We apply the smoothness loss on the rendered albedo, metallic and roughness map. Furthermore, We assume a near-natural white incident light to obtain a plausible lighting materials disentangle. We therefore apply the white light regularizations on the environment map as:
$L_{\text{white}}=\sum_{\text{channel}}(\mathbf{L}_{\text{channel}}-\frac{1}{3}\sum_{\text{channel}}\mathbf{L}_{\text{channel}}),{\text{channel}}\in\{R,G,B\}$ (23)
@@ 3.5 Novel pose animation under novel illuminations
After optimization, we can get the canonical PBR-aware 2DGS $\mathbf{G}$. To achieve novel pose animation under novel illuminations, we first deform the optimized canonical PBR-aware 2DGS to the observation space with the given novel pose which is unseen in the training process following the equation 7. We then use the environment map of the novel illuminations to build a cubemap. For each viewpoint, we query the cubemap to calculate the PBR color for each Gaussian at observation space with the equation 13. Finally, we apply the splatting process (equation 14) to render the avatar at given novel illuminations.
@@ 4.1 Evaluation datasets
To validate the effectiveness of our proposed methods, we use a synthetic dataset (RANA [92]) and two real-world datasets (PeopleSnapshot [104], ZJU-MoCap [38, 105]) for performance evaluation.
RANA [92]. RANA [92] is a synthetic dataset that includes the ground-truth albedo, normal, and relighting results at novel pose for evaluation. The training view is a subject with an A-pose rotating in front of a camera in the scene. The testing set includes the same subject under unknown illumination with random poses. To quantitatively evaluate the physical material of the reconstructed human, we use 8 subjects from the RANA dataset [92] by following the same protocols in IntrinsicAvatar [29].
PeopleSnapshot [104] PeopleSnapshot captures subjects that hold a nearly A-pose and rotate in front of the camera. The subjects are under natural illumination from the real world. We follow previous works [40, 35] and select 4 subjects from the dataset for qualitative evaluation. After optimization, our reconstructed avatar can be rendered in a novel pose under novel illuminations.
ZJU-MoCap [38, 105] In ZJU-MoCap, the subjects captured in the video may conduct very complex motions. The shadow effect during the complex motion challenges our physically based rendering. We use the dataset to show that our method can achieve human intrinsic disentangles from a monocular video with complex human motions.
@@ 4.3 Evaluation Metrics
Figure 4: Qualitative Results on the RANA Dataset. We visualize the PBR image, albedo, and normal from the training view. We then relight the clothed human avatars at 2 different novel poses under the novel illuminations.
@@
  Method Relighting (novel pose) Normal Error $\downarrow$ Albedo Training Time
@@
Albedo: PSNR/SSIM/LPIPS: We use the image quality metrics to evaluate the albedo estimation from the training view. Since ambiguity exists between the albedo scale and light intensity, we follow previous works [22, 29] to align the estimated albedo with a scale of the ground-truth albedo.
Normal Estimation Error: The metric evaluates the normal errors between the predicted normal and the ground-truth normal from the training view. We empirically find that the rendered normal is better.
@@ 4.4 Experimental Results
Figure 6: Qualitative results on the ZJU-MoCap [38] dataset. In the left 6 columns, we show the rendered image, estimated albedo, roughness, metallic, normal, and the approximated occlusion. In the right 3 columns, we visualize the relighting results under different novel illuminations at different poses.
@@
Training details. We first train 5,000 steps for the image reconstruction stage. We apply the normal-depth consistency regularizer after 3,000 steps, from which we can start to align the splat normal to the optimized geometry. We then conduct 3,000 steps to optimize the PBR materials. The overall training step is 8,000 steps, which consumes about 40 minutes on a single Nvidia 3090Ti GPU (with peak 13GB GPU memory usage), which is almost 6 times faster than IntrinsicAvatar. We set the $\lambda_{\text{NC}}=0.05$ for the normal-consistency loss. The weights in the smooth loss ($\lambda_{\mathbf{a}},\lambda_{r},\lambda_{m}$) are set to be 0.02. The weight of $\lambda_{\text{white}}$ is set to be 0.1. We follow the learning rate in the original Gaussian attribute in Gaussian Splatting [32, 33]. For the additional parameters of the PBR attribute and the learnable environment light, we set the learning rate the same as that of the spherical harmonics appearance. In the image reconstruction stage, the overall loss is eq. 15, while in the PBR optimization stage, the overall loss is eq. 20.
@@ 4.4.1 Comparison on synthetic dataset
We compare with the baselines in Tab.I 333 We re-run the baselines [25, 29] with the official implementation provided by the authors:
@@
and Fig.4. We can observe that R4D* [25] cannot reconstruct the surface geometry accurately. Since such a method relies on the time-conditioned embedding during optimization, thus, it cannot generalize to the novel poses. Moreover, its implicit representation makes it lack of the details. While Intrinsic Avatar [29] can generalize to the novel pose and its multi-level Hash-Grids [31] representation provide much more details, the albedo details (in the man’s face on the left of Fig. 5) are still missing. For the female on the right of Fig. 5, the dress surface reconstructed by  [29] tends to overfit the texture, while our methods can reconstruct the smooth surface for the dress. Furthermore, our methods can bring high light during relighting, which brings superiority for relighting (novel pose). The quantitative results in Tab. I also demonstrate this. Compared with R4D and R4D*, our methods reduce $75.8\%$ and $62.6\%$ normal errors and significantly improve the PSNR of estimated albedo by $19.2\%$. Compared with IntrinsicAvatar, our methods achieve comparable estimated albedo quality (a lower PSNR (0.46 dB), but a higher SSIM and LPIPS) and more accurate relighting quality (a higher PSNR (0.62 dB)). Further, compared with SOTA methods, our methods can achieve a 5×5$\times$5 × faster in training speed and a 100×100$\times$100 × faster in rendering speed.

### v3

@@ 3.4 Reconstruct animatable PBR-aware 2DGS from monocular videos
Our training process consists of two stages. In the first stage, we apply an Image Reconstruction loss to capture the rough shape and appearance of the dynamic humans. This allows us to obtain an initial estimate of the avatar’s geometry and appearance. In the second stage, we use PBR techniques to further refine the avatar’s material properties and achieve accurate surface reconstruction. Specifically, we optimize the model parameters $\mathbf{G},\widehat{\mathbf{W}},\widetilde{\mathbf{B}},\widetilde{\mathbf{W}}$ using the ground-truth images $\mathbf{I}_{1},...,\mathbf{I}_{M}$} and the corresponding human poses $\boldsymbol{\theta}_{1},...,\boldsymbol{\theta}_{M}$}
Training Stage I – Image Reconstruction. In the first stage, we aim to reconstruct the rough shape for clothed avatars with view-dependent color represented as spherical harmonics. We apply the image reconstruction loss and a regularization term on these Gaussian parameters, and the normal-depth consistency loss proposed by [33]. Denote the rendered image as $\hat{\mathbf{I}}$, the training loss at this stage is:
$L_{\text{RF}}=\ell_{1}(\hat{\mathbf{I}},\mathbf{I})+\lambda_{\text{SSIM}}L_{\text{SSIM}}(\hat{\mathbf{I}},\mathbf{I})+L_{\text{reg}}+\lambda_{\text{NC}}L_{\text{NC}}\text{.}$ (15)
We only render the foreground avatar and apply the $L_{\text{RF}}$ loss on it, without modeling the background. Due to the discrete Gaussian representation, the reconstructed cloth avatars may overfit to the sparse 2D observations without regularization. Following prior work  [39], we regularize the standard deviation $\sigma_{k}$ of the Gaussian attributes within the KNN neighborhood of each Gaussian $\mathbf{G}_{k}$. This leads to:
$L^{\sigma}_{k}=\sum_{\text{attr}_{k}\in\{\mathbf{R}_{k},\mathbf{S}_{k},o_{k},\mathbf{c}_{k},\Delta\mathbf{W}_{k}\}}\lambda_{\text{attr}_{k}}\sigma(\text{attr}_{k})\text{,}$ (16)
where $\sigma_{\text{attr}_{k}}$ denotes the standrd variation of the Gaussian attributes in the KNN neighborhood. Furthermore, we encourage learnable blending weights, and the latent bones for cloth animation for each Gaussian to be small, thus alleviating the overfitting for the cloth deformation. We follow the previous work [35] to store the learnable blending weights in a 3D voxel and apply a tri-linear interpolation to query the blending weight $\Delta W_{k}$ for each Gaussian. We therefore regularize the learnable blending weights to be small for both the SMPL bones and latent bones with:
$L_{k}^{\text{norm}}=\lambda_{\widehat{\mathbf{W}}}\|\Delta\mathbf{W}_{k}\|_{2}+\lambda_{\widetilde{\mathbf{W}}}\|\widetilde{\mathbf{W}}_{k}\|_{2}\text{.}$ (17)
@@
$L_{\text{reg}}=\frac{1}{N_{\text{gs}}}\sum_{k=1}^{N_{\text{gs}}}L_{k}^{\sigma}+L_{k}^{\text{norm}}\text{.}$ (18)
To ensure the orientation of each 2D Gaussian align with the actual surface normal and suitable for physically based rendering. We follow 2DGS [33] and align the splat’s normal with the gradients of the depth maps as follows:
$L_{\text{NC}}=\sum_{i}w_{i}(1-\mathbf{n}_{i}^{T}\mathbf{N})\text{,}$ (19)
where $w_{i}=o_{i}\mathcal{G}_{i}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-o_{j}\mathcal{G}_{j}(\mathbf{u}(x)))$ denotes the blending weight of the $i$-th intersection, $\mathbf{n}_{i}$ is the normal of the splat facing the camera, $\mathbf{N}$ is normal derived from depth map, Compared with the $\ell_{1}$ loss utilized in prior works [46, 48], our normal-consistency loss yields smoother surfaces with more consistent Gaussian orientation along the ray, as shown in our ablation study.
Training Stage II – PBR optimization. At the PBR stage, we first apply the image reconstruction loss on the PBR image. Denote the rendered image of PBR as $\mathbf{\hat{I}}_{\text{pbr}}(\mathbf{a},m,r,\mathbf{n})$, which is computed from albedo, materials, and splat normal. The overall loss at the PBR stage is:
@@
$\lambda_{\text{NC}}L_{\text{NC}^{*}}+\lambda_{\text{white}}L_{\text{white}}\text{,}$
where the $L_{\text{NC}^{*}}$ is slightly different from the $L_{\text{NC}}$ used in the previous stage. With the original normal-consistency loss from the previous stage, we can not reconstruct totally accurate geometry, especially under the extreme high light conditions as shown in Fig 3 and the original 2DGS [33] paper. At the PBR stage, the Gaussian appearance correlates with the splat normal. Thus, we can improve geometry during optimization. Different with GS-IR [46], we cannot freeze Gaussian shapes. However, directly optimizing Eq. 19 leads to inconsistent splat normals due to inaccurate geometry. Discarding the normal consistency regularization makes splat normals erratic Fig. 11. Therefore, we propose a progressive optimization strategy for accurate geometry and consistent splat normals:
@@
$+\lambda_{\text{NC}_{2}}\sum_{i}w_{i}(1-\mathbf{n}_{i}^{T}\mathbf{N})\text{,}$
where $\text{sg}(\cdot)$ denotes stop gradient. At the initial stage, we set $\lambda_{\text{NC}_{2}}=0,\text{and}\,\lambda_{\text{NC}_{1}}=1$, which means we utilize the splat normal to optimize the geometry $\mathbf{N}$. Then we set $\lambda_{\text{NC}_{1}}=0,\text{and}\,\lambda_{\text{NC}_{2}}=1$ to force the splat normal to be aligned with the fixed geometry. We visualize the splat normal and depth derived normal during the optimization stage at Fig 3. To overcome the ambiguity between the lighting and PBR materials and obtain a plausible solution for PBR materials, we follow  [47] to utilize a smoothness loss which forces the materials and albedo to be smooth at the planar surfaces, where the image gradients are absent. This can be written as:
@@
$+\frac{1}{N_{\text{pixel}}}\sum_{i,j}|\partial_{\mathbf{y}}\mathbf{a}_{ij}|e^{-\|\partial_{\mathbf{y}}\mathbf{I}_{ij}\|}\text{,}$
where $N_{\text{pixel}}$ denotes the numbe of pixels of the image. I represents ground-truth images, $\partial_{\mathbf{x}},\partial_{\mathbf{y}}$ represents the gradients of an image or material map in $\mathbf{x}$ and $\mathbf{y}$ directions. We apply the smoothness loss on the rendered albedo, metallic and roughness map. Furthermore, we assume a near-natural white incident light to obtain a plausible disentanglement of lighting and materials. We therefore apply the white light regularization on the environment map as:
$L_{\text{white}}=\sum_{\text{channel}}(\mathbf{L}_{\text{channel}}-\frac{1}{3}\sum_{\text{channel}}\mathbf{L}_{\text{channel}}),{\text{channel}}\in\{R,G,B\}\text{.}$ (23)
@@ 3.5 Novel pose animation under novel illuminations
After optimization, we can get the canonical PBR-aware 2DGS $\mathbf{G}$. To achieve novel pose animations under novel illumination, we first deform the optimized canonical PBR-aware 2DGS to the observation space with the given novel pose which is unseen in the training process following the equation 7. We then use the environment map of the novel illuminations to build a cubemap. For each viewpoint, we query the cubemap to calculate the PBR color for each Gaussian at observation space with the equation 13. Finally, we apply the splatting process (equation 14) to render the avatar at given novel illumination.
@@ 4.1 Evaluation datasets
To validate the effectiveness of our proposed method, we use a synthetic dataset (RANA [97]) and two real-world datasets (PeopleSnapshot [109], ZJU-MoCap [80, 110]) for performance evaluation.
RANA [97]. RANA [97] is a synthetic dataset that includes ground-truth albedo, normal, and relighting results at novel pose for evaluation. Each training view is a subject with an A-pose rotating in front of a camera in the scene. The testing set includes the same subject under unknown illumination with random poses. To quantitatively evaluate the physical materials of the reconstructed human, we use 8 subjects from the RANA dataset [97] following the same protocols in IntrinsicAvatar [29].
PeopleSnapshot [109] PeopleSnapshot captures subjects that hold a nearly A-pose and rotate in front of the camera. The subjects are under natural illumination in the real world. We follow previous works [39, 35] and select 4 subjects from the dataset for qualitative evaluation. After optimization, our reconstructed avatar can be rendered in a novel pose under novel illuminations.
ZJU-MoCap [80, 110] In ZJU-MoCap, the subjects captured in the video conduct very complex motions. The shadow effect during the complex motion challenges our physically based rendering. We use the dataset to show that our method can achieve human intrinsic disentangles from a monocular video with complex human motions.
@@ 4.3 Evaluation Metrics
Figure 4: Qualitative results on the RANA dataset. We visualize the PBR image, albedo, and normal from the training view. We then relight the clothed human avatars at 2 different novel poses under the novel illuminations. As highlighted by the red rectangles, IntrinsicsAvatar [29] may optimize the fake normal to fit the appearance, while our method provides a more realistic normal for better relighting.
@@
  Method Relighting (novel pose)
Normal
Error
$\downarrow$
Albedo Training Time
@@
Albedo: PSNR/SSIM/LPIPS: We use the image quality metrics to evaluate the albedo estimation from the training view. Since ambiguity exists between the albedo scale and light intensity, we follow previous works [22, 29] to align the estimated albedo scale to the ground-truth albedo scale.
Normal Estimation Error: The metric evaluates the normal errors between the predicted normal and the ground-truth normal from the training view.
@@ 4.4 Experimental Results
Figure 6: Qualitative results on the ZJU-MoCap [80] dataset. In the left 6 columns, we show the rendered image, estimated albedo, roughness, metallic, normal, and the approximated occlusion. In the right 3 columns, we visualize the relighting results under different novel illuminations at different poses. Top row: for the avatar with no complex wrinkles on the cloth, our method reconstruct smooth and reasonable physical propertis. Bottom row: for the avatar with complex wrinkles on the cloth, although our method can still work well on most regions, it cannot calculate the physical properties correctly at the winkle area as highlighted by the red rectangles.
@@
Training details. The image reconstruction stage includes 5,000 steps. We apply the normal-depth consistency regularizer after 3,000 steps, from which we can start to align the splat normal to the optimized geometry. We then conduct 3,000 steps to optimize the PBR materials. The overall training step is 8,000 steps, which consumes about 40 minutes on a single Nvidia 3090Ti GPU (with peak 13GB GPU memory usage), which is almost 6 times faster than IntrinsicAvatar. We set the $\lambda_{\text{NC}}=0.05$ for the normal-consistency loss. The weights in the smooth loss ($\lambda_{\mathbf{a}},\lambda_{r},\lambda_{m}$) are set to be 0.02. The weight of $\lambda_{\text{white}}$ is set to be 0.1. We follow the learning rate in the original Gaussian attribute in Gaussian Splatting [32, 33]. For the additional parameters of the PBR attribute and the learnable environment light, we set the learning rate the same as that of the spherical harmonics appearance. In the image reconstruction stage, the overall loss is eq. 15, while in the PBR optimization stage, the overall loss is eq. 20.
@@ 4.4.1 Comparison on synthetic dataset
We compare with the baselines in Tab.I  [^3] We re-run the baselines [25, 29] with the official implementation provided by the authors:
@@
and Fig.4. We can observe that R4D* [25] cannot reconstruct the surface geometry accurately. Since such a method relies on the time-conditioned embedding during optimization, it cannot generalize to the novel poses. Moreover, its implicit representation lacks detail. While Intrinsic Avatar [29] can generalize to the novel pose and its multi-level Hash-Grids [31] representation provide much more details, the albedo details (in the man’s face on the left of Fig. 5) are still missing. For the female on the right of Fig. 5, the dress surface reconstructed by  [29] tends to overfit the texture, while our methods can reconstruct the smooth surface for the dress. Furthermore, our method can bring high light during relighting, which brings superiority for relighting (novel pose) as shown in the relighting results on the back of the female in the Fig. 4. The quantitative results in Tab. I also demonstrate this. Compared with R4D and R4D*, our methods reduce $75.8\%$ and $62.6\%$ normal errors and significantly improve the PSNR of estimated albedo by $19.2\%$. Compared with IntrinsicAvatar, our methods achieve comparable estimated albedo quality (a lower PSNR (0.46 dB), but a higher SSIM and LPIPS) and more accurate relighting quality (a higher PSNR (0.62 dB)). Further, compared with SOTA methods, our methods can achieve a $5\times$ faster in training speed and a $100\times$ faster in rendering speed.